OpenStudy (some.random.cool.kid):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (some.random.cool.kid):

can anyone help

OpenStudy (owlcoffee):

First off, try solving it and then tweak the variables a little with some algebraic properties so we get a new, but equivalent system of linear equations.

OpenStudy (some.random.cool.kid):

how do i solve this

OpenStudy (some.random.cool.kid):

ill come back when your finished writing brb

OpenStudy (owlcoffee):

Okay, let's first look at the system of equation, I'll take the 1st equation and solve for one variable: \[6x+2y=-6\] \[2y=-6-6x\] \[y=\frac{ -6x-6 }{ 2 }\] \[y=\frac{ 6(-x-1) }{ 2 }=3(-x-1)=-3x-3\] \[y=-3x-3\] After solving for the variable I chose, I'll replace it on the 2nd equation: \[3x-4y=-18\] \[y=-3x-3\] then: \[3x-4(-3x-3)=-18\] \[3x+12x+12=-18\] \[15x=-30\] \[x=\frac{ -30 }{ 15 }\] \[x=-2\] having found "x", all we have to do is replace it on the "y" value we had found on the first equation: \[y=-3x-3\] \[x=-2\] then: \[y=-3(-2)-3\] \[y=6-3=2\] \[y=2\] and thus, solving the problem.

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

can i show you the list of answers though

OpenStudy (some.random.cool.kid):

here look

OpenStudy (some.random.cool.kid):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 ANSWERS BELOW!!! ___________________ A. 12x + 4y = -12 15x = -30 B. 8x + 4y = -4 14x = -10 C. 6x - 8y = -36 -6y = -42 D. 6x - y = -15 3y = 9

OpenStudy (some.random.cool.kid):

there

OpenStudy (some.random.cool.kid):

@ganeshie8

OpenStudy (some.random.cool.kid):

how to solve

OpenStudy (some.random.cool.kid):

@ganeshie8 you there

ganeshie8 (ganeshie8):

6x + 2y = -6 3x - 4y = -18

OpenStudy (some.random.cool.kid):

what

ganeshie8 (ganeshie8):

multiply first equation by 2 you get 12x + 4y = -12 3x - 4y = -18 yes ?

OpenStudy (some.random.cool.kid):

yeah

ganeshie8 (ganeshie8):

next add both the equations vertically

OpenStudy (some.random.cool.kid):

ok

ganeshie8 (ganeshie8):

do it

OpenStudy (some.random.cool.kid):

ok wait how exatly

OpenStudy (some.random.cool.kid):

ok nvm then.

ganeshie8 (ganeshie8):

12x + 4y = -12 3x - 4y = -18 -------------------- 15x + 0 = -30

OpenStudy (some.random.cool.kid):

oh

OpenStudy (some.random.cool.kid):

ok

ganeshie8 (ganeshie8):

you just add them column by column like above ^

OpenStudy (some.random.cool.kid):

ok thanks i get it now

ganeshie8 (ganeshie8):

you're welcome!

OpenStudy (owlcoffee):

oh, Sorry if I didin't response, I lost connection X.x

OpenStudy (some.random.cool.kid):

oh

OpenStudy (some.random.cool.kid):

anyways bye guys im going to close the question now